OpenStudy (anonymous):

. Suppose that A = D + E and that D/A = 0.75. Solve for D/E.

OpenStudy (anonymous):

\[ \frac D A = \frac 3 4\\ A =\frac 4 3 D\\ \frac 4 3 D = D +E\\ \frac 4 3 \frac D E= \frac D E +1 \\ \] Can you finish it?

OpenStudy (anonymous):

\[ \frac D A = \frac 3 4\\ A =\frac 4 3 D\\ \frac 4 3 D = D +E\\ \frac 4 3 \frac D E-\frac D E=1 \\ \frac 1 3 \frac D E =1 \\ \frac D E=3 \]

OpenStudy (anonymous):

I'm going to try give me a second

OpenStudy (anonymous):

Try to understand what I posted above.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thank you